OpenStudy (anonymous):

A company that manufactures bicycles has a fixed cost of $100,000. It costs $100 to produce each bicycle. The selling price per bike is $300. Write the cost function, C. Write the revenue function, R. Determine the break-even point. Describe what this means. Hint: In solving this exercise, let x represent the number of bicycles produced and sold.

OpenStudy (saifoo.khan):

online college?

OpenStudy (anonymous):

yeah playing catch up... fel behind a bit

OpenStudy (saifoo.khan):

Fixed cost = $100,000 -- this is the cost the money will spend even if they produce no bikes. Cost per bike = $100 -- this is an operational cost. It costs $100 to build one bike. Cost of x bikes = x*$100 Revenue per bike = $300 -- this is the amount of money that will be made when each bike is sold. It's the price of the bike. Revenue for x bikes = x*$300 x = number of bikes produced and sold i. C = cost = fixed cost + variable cost C = $100,000 + $100*x ii. R = revenue = amount of money made R = x*$300 iii. The break-even point is the point at which your Total Cost (C) is the same as your Total Revenue (R). In other words, it's the point at which you make $0. C = R $100,000 + $100*x = $300*x $100,000 = $200*x x = 500 After building 500 bicycles you break even.

OpenStudy (saifoo.khan):

Oh cool. its good to study! :D

OpenStudy (anonymous):

thnks

OpenStudy (saifoo.khan):

welcome. please hit good answer.

OpenStudy (saifoo.khan):

\begin{array}l\color{red}{\text{t}}\color{orange}{\text{h}}\color{#9c9a2e}{\text{n}}\color{green}{\text{k}}\color{blue}{\text{s}}\color{purple}{\text{!}}\color{purple}{\text{}}\end{array}

OpenStudy (anonymous):

how are you with graphs.